# Forecasting Automotive Supply Chain Disruption with Heterogeneous Time Series

### v1

@@ 3 Data Description and Feature Engineering
Table 1: Full List of predictive features and their brief descriptions. The features are selected the reflect of aspects of Inventory, Capacity, Utilization and Processing Time in the classical Factory Physics.
In this section, we first describe how we construct and gather the data necessary for tracking suppliers’ delivery performance and eventually forecasting time-to-disruption events. Ford contracts with many suppliers to ship critical vehicle parts to our manufacturing plants. Each unique part ID is assigned a ”will-make” date, indicating the scheduled date for its use in the vehicle assembly line. Disruption events occur when a part’s promised delivery date exceeds its designated will-make date.
@@
A pivotal result in classical manufacturing management (see hopp2011factory) is the understanding that, in the presence of variability, three key factors are essential for synchronizing demand and production at lower costs and higher service levels: capacity, inventory, and cycle time. This concept is technically articulated through Little’s Law and Kingman’s equation. According to Little’s Law, inventory ($L$), measured in quantity units, is proportional to throughput ($\lambda$), in quantity units per time, and cycle time ($W$): $L=\lambda\cdot W$. Kingman’s equation, also known as the VUT equation, provides the expected waiting time or cycle time ($W$) as a function of variability, utilization, and process time: $\mathbb{E}[W]=V\cdot U\cdot T$ (as discussed in hopp2011factory). This equation was developed by the British mathematician John Kingman in $1961$ using queuing theory (see kingman1961). The first factor, variability, captures the randomness of arrival and service times; the second factor, utilization, reflects the usage of the workstation or assembly line; and the third factor, process time, represents the average processing time of the assembly line.
In this paper, instead of explicitly spelling out the equation for time-to-disruption, we rely on historical trends and patterns to develop a statistical model to predict disruption times, which will be discussed in Section 4. We recognize that many principles from Factory Physics still apply in the modern era. Consequently, we construct our dataset with features related to capacity, inventory, utilization, and processing time.
Our dataset comprises many time series collected from January $1$, $2021$, to December $31$, $2023$. Each time series corresponds to a combination of supplier, Ford plant, and vehicle part IDs, representing an agreement between a supplier and a Ford plant to deliver vehicle parts. These multivariate time series consist of $21$ features recorded daily, chosen to reflect the manufacturing process’s capacity, inventory, utilization, and processing time aspects. Table 1 lists the complete features used in this study. We also track the time-to-disruption at each time point, with the exact time in days until a disruption event. The time-to-events are known since they are historical events.
@@
For each row feature, the response variable is the time in days until a disruption event from the last day of the $28$-day sequence—these times until disruption are truncated to lie between 0 and 366 days. The value 366 serves as a unique token indicating that the time series is predicted not to experience any disruptions within the annual fiscal year.
Notationally, our data consists of $(d_{i},X_{i},t_{i},y_{i}),\;i=1,2,\ldots,n$, where $X_{i}$ represents the $28$-day sequence features/covariates, $t_{i}$ is the time in days until disruption, $y_{i}$ is an indicator with one signifying that the time series resulted in disruption and 0 otherwise, and $d_{i}$ is a tuple containing three identifiers for the site, Ford plant, and vehicle part. Figure 1 illustrates the Pearson correlation among $21$ features, revealing no pairs of features with perfect correlation. Furthermore, Figure 2 demonstrates that roughly $25$% of all the time series are right-censored.
@@ 4 Methodology
In the preceding section, we described the construction of our dataset through the lenses of capacity, inventory, utilization, and processing time, as outlined in Factory Physics. This section details our methodology for modeling this dataset to predict the time until disruption.
Given a sample $(d_{1},X_{1},t_{1},y_{1}),(d_{2},X_{2},t_{2},y_{2}),\ldots,(d_{n},X_{n},t_{n},y_{n})$, we model the times until disruption using a Survival Analysis model, as described in Section 2.1. Specifically, our chosen hazard function has a logistic form. The function $\phi_{k}(\cdot),k=1,2,\ldots,366$ is a smooth, nonlinear function that models the relationship between the feature $X_{i}$ and the time until disruption $t_{i}$.
$h(\tau_{k}|X_{i})=\frac{1}{1+exp(-\phi(X_{i}))},k=1,2,\ldots,366$ (5)
We estimate the functions $\phi_{1},\ldots,\phi_{366}$ using a deep learning neural network based on the Attention Seq2Seq with bidirectional GRUs described in Section 2.3. In this network, the feature $X_{i}$ represents a sequence of 21 metrics collected over 28 days, as detailed in Section 3. This sequence is fed into the Encoder, which transforms it into 28 hidden states and a context vector—a weighted average of these hidden states with Attention Mechanism. The Encoder then processes these hidden states and the context vector to generate 366 elements, $\phi_{1},\phi_{2},\ldots,\phi_{366}$. These values, ${\phi_{k}}_{k=1}^{366}$, are used to compute the hazard functions, which are crucial to the final loss function employed to train the neural network via backpropagation. The loss function is the negative log-likelihood, given by:
@@
We propose a important change to the Seq2Surv model (li2022), specifically enhancing the Encoder Network to better model heterogeneity among plants, sites, and parts. By incorporating Neural Embeddings for plant, site, and part IDs, we aim to effectively capture the dataset’s diversity. We refer to our modified model as Heterogeneous Sequence-to-Survival.
As illustrated in Figure 3 each site ID, plant ID, and vehicle part ID is transformed into vector embeddings using neural network embedding layers. These embeddings are then concatenated and passed through two fully connected neural network layers, which generate non-linear group effect signals. These signals are subsequently added to the hidden states of the bi-directional GRUs and concatenated with the context vector. This methodological adjustment is crucial for accounting for the heterogeneity in the dataset. Empirical observations indicate that the disruption times for different combinations of sites, plants, and parts can exhibit distinct and sometimes contradictory behaviors, highlighting the need for our approach to capture these nuanced variations. The Encoder encapsulates the essential representation for each $28$-day sequence feature $X_{i}$. Intuitively, this representation consists of two components: an individual-specific component and a group component related to the combination of site, plant, and part values. The individual-specific information is directly inferred from the feature values of $X_{i}$, ensuring that the unique characteristics of each sequence are accurately modeled. Meanwhile, the group component is derived from the site, plant, and part values, allowing the model to incorporate the distinct group behaviors from the data. The individual representation captures the idiosyncrasies of each sequence, while the group representation integrates the broader, systematic effects associated with different combinations of site, plant, and part values. This modified architecture enhances the model’s ability to predict disruption times with greater accuracy and reliability, accommodating the inherent variability within the dataset.
@@ 5.1 Predictive Performance
Kansas City Assembly Plant 0.85 0.82 0.23 0.05 0.04 0.68
Michigan Assembly Plant 0.95 0.9 0.22 0.02 0.00 0.75
Dearborn Truck Plant 0.95 0.81 0.15 0.04 0.01 0.8
Ohio Assembly Plant 0.95 0.86 0.21 0.03 0.01 0.75
Kentucky Truck Plant 0.85 0.81 0.25 0.06 0.04 065
@@
In this section, we detail the approach for evaluating the proposed model at Ford to ensure its quality. The trained model estimates the time until a disruption event occurs related to the shipment of vehicle parts between a supplier and a Ford plant. To assess the model’s performance, we modify and adapt the standard classification metrics to fit our specific needs.
In binary classification, the two standard metrics for evaluation are Precision and Recall (see buckland1994). Binary classification involves two labels: positive and negative. For a given classifier, True Positives (TP) are the instances where the classifier correctly identifies positives. True Negatives (TN) are the instances where the classifier correctly identifies negatives. False Positives are instances where the classifier incorrectly predicts positives when the true label is negative. Conversely, False Negatives (FN) are instances where the classifier incorrectly predicts negatives when the true label is positive. Formally, Precision and Recall are defined as,
@@
Precision and recall range from 00 to $1$. The closer these values are to $1$, the better the classifier’s performance. Specifically, recall quantifies the proportion of actual positive data points that the classifier successfully identifies. Precision, on the other hand, measures the proportion of predicted positive cases that are, in fact, positive. Informally, recall measures how effectively a classifier identifies positive events (disruptions in this context), while precision measures the frequency with which the classifier avoids generating false alarms. These two metrics offer a comprehensive view of a model’s classification performance.
Given that the model proposed in Section 4 estimates the time until a disruption, we first define a forecasting horizon, $\Delta$ (in days), and a margin-of-error, $\epsilon$ (in days), to evaluate its performance. We introduce the following definitions to adapt the concepts of True Positives, False Positives, True Negatives, and False Negatives to our specific problem. Given an observation time $t_{0}$, and the forecasting window $\Delta$ and margin-of-error $\epsilon$, $t_{0}+\Delta$ represents a time point $\Delta$ days from $t_{0}$. Adapted True Positives (ATP) are defined as the number of disruptions that occur exactly $\Delta$ days from $t_{0}$ and are estimated by the model within $\epsilon$ days of $t_{0}+\Delta$. The difference between the model’s estimated time and $t_{0}+\Delta$ must fall within the range $[-\epsilon,\epsilon]$ days. Adapted False Positives (AFP) are instances where the model predicts a disruption time exactly $\Delta$ days from $t_{0}$, but no disruptions actually occur within $\epsilon$ days of $t_{0}+\Delta$.
Conversely, Adapted True Negatives (ATN) are instances where the model estimates the disruption time to be more than $\Delta$ days from $t_{0}$, and no disruptions occur between $t_{0}$ and $t_{0}+\Delta$. Adapted False Negatives (AFN) are instances where the model predicts a disruption time beyond $\Delta$ from $t_{0}$, while disruptions actually occur within $\Delta$ days from $t_{0}$.
With these definitions, Adapted Recall is computed as the ratio of Adapted True Positives (ATP) to the total of ATP and Adapted False Negatives (AFN). Likewise, Adapted Precision is the proportion of ATP relative to the sum of ATP and Adapted False Positives (AFP). Additionally, we provide the Normalized Confusion Matrix in Table 2 also known as the error matrix. This table offers a detailed visualization of the four cases: True Positives, False Negatives, False Positives, and True Negatives. The normalization is done by dividing each case by the total number of cases.
@@
Seq2Surv on Kansas Assembly Plant Data 0.62 0.3
Seq2Surv on Kentucky Truck Plant 0.52 0.21
Heterogeneous Seq2Surv on Kansas Assembly Plant Data 0.85 0.82
Heterogeneous Seq2Surv on Kentucky Truck Plant 0.85 0.81
Table 3: Model Performance Comparison of Seq2Surv and Heterogeneous Seq2Surv on Kansas Assembly Plant and Kentucky Truck Plant Datasets. Heterogeneity presents a critical challenge for Seq2Surv.
We conducted 20 iterations of Quality Assurance (QA) testing from August 1, 2023, to January 13, 2024 for the result in Table 2. For each observation week, we trained the model on the cumulative data up to that week and used it to predict disruptions in the near future. We computed Adapted Precision, Adapted Recall, and the Normalized Confusion Matrix using definitions for Adapted True Positives, True Negatives, False Positives, and False Negatives, with parameters forecasting horizon $\Delta=28$ days and margin-of-error $\epsilon=7$ days. The metrics in Table 2 are averaged over the 20 iterations. The Quality Assurance (QA) evaluation was conducted at five selected Ford manufacturing facilities in North America: Kansas City Assembly Plant, Michigan Assembly Plant, Dearborn Truck Plant, Ohio Assembly Plant, and Kentucky Truck Plant.
The model demonstrated on average precision exceeding $0.85$ and a recall surpassing $0.8$ across the five plants during the 20-week QA period. Additionally, the normalized confusion matrix indicates that the combined error rate of false positive and false negative errors was below $10$% on average. Furthermore, we present performance comparison between Seq2Surv, as described in li2022, and our Heterogeneous Seq2Surv proposed in Section 4, as shown in Table 3. The metrics provided are the averages obtained from 20 iterations of QA testing. Our analysis reveals that the heterogeneity in plant, supplier, and part groups poses significant challenges for Seq2Surv. However, by incorporating heterogeneous behaviors in the data into the model, our Heterogeneous Seq2Surv demonstrates substantially improved performance. For a comparison between Seq2Surv and other models, please refer to li2022.
@@ 6 Conclusion
In this paper, we first outline the process of constructing a dataset of multivariate time series for forecasting first-tier supply chain disruptions. We carefully selected features representing key aspects of classical Factory Physics, such as capacity, inventory, utilization, and processing.
The dataset, comprising over five hundred thousand individual time series, presents significant technical challenges due to its complexity and scale. Although these time series exhibit some commonalities, they also demonstrate substantial heterogeneity within specific subgroups, making traditional industrial and statistical models inadequate for capturing the intricate dynamics and managing the extensive data volume. To address these challenges, we propose a novel methodology that integrates an enhanced Attention Sequence-to-Sequence Deep Learning architecture with Neural Network Embeddings to model group effects, combined with a Survival Analysis model. This approach is designed to effectively capture complex heterogeneous data patterns related to operational disruptions. Our model has achieved strong performance, with at least $85$% precision and $80$% recall during the Quality Assurance (QA) phase across Ford’s five North American plants.
To mitigate the common critique of Machine Learning models as ”black boxes,” we employed the SHAP (SHapley Additive exPlanations) framework to elucidate feature importance in the model’s predictions. This technique provides valuable insights that can guide actionable strategies for business partners. Our work highlights the potential of advanced machine learning techniques to manage and mitigate supply chain risks effectively within the automotive industry.

### v3

@@ 3 Data Description and Feature Engineering
Table 1: Full List of predictive features and their brief descriptions. The features are selected the reflect of aspects of Inventory, Capacity, Utilization and Processing Time.
In this section, we first describe how we construct and gather the data necessary for tracking suppliers’ delivery performance and eventually forecasting time-to-shortfall events. Ford contracts with many suppliers to ship critical vehicle parts to our manufacturing plants. Each unique part ID is assigned a “will-make” date, indicating the scheduled date for its use in the vehicle assembly line. Shortfall events occur when a part’s promised delivery date exceeds its designated will-make date.
@@
A pivotal result in classical manufacturing management is the understanding that, in the presence of variability, three key factors are essential for synchronizing demand and production at lower costs and higher service levels: capacity, inventory, and cycle time (hopp2011factory). This concept is technically articulated through Little’s Law and Kingman’s equation (see hopp2011factory). According to Little’s Law, inventory ($L$), measured in quantity units, is proportional to throughput ($\lambda$), in quantity units per time, and cycle time ($W$): $L=\lambda\cdot W$. Kingman’s equation, also known as the VUT equation, provides the expected waiting time or cycle time ($W$) as a function of variability, utilization, and process time: $\mathbb{E}[W]=V\cdot U\cdot T$ (as discussed in hopp2011factory). This equation was developed by the British mathematician John Kingman in $1961$ using queuing theory (see kingman1961). The first factor, variability, captures the randomness of arrival and service times; the second factor, utilization, reflects the usage of the workstation or assembly line; and the third factor, process time, represents the average processing time of the assembly line.
In practice, accurately estimating variability, capacity, and utilization can be difficult. To address this, we use Statistical Machine Learning to build a model that learns from historical trends and patterns to predict shortfalls, as discussed in Section 4. For the model to be effective, the training dataset is designed with data features strongly correlated with variability, capacity, utilization, and processing time. It comprises time series collected from January $1$, $2021$, to December $31$, $2023$. Each time series corresponds to a combination of supplier, Ford plant, and vehicle part IDs, representing an agreement between a supplier and a Ford plant to deliver vehicle parts. Table 1 lists the complete features used in this study. We also track the time-to-shortfall at each time point, with the exact time in days until a shortfall event. The time-to-events are known since they are historical events.
@@
For each row feature, the response variable is the time in days until a shortfall event from the last day of the $28$-day sequence— these times until shortfalls are truncated to lie between 0 and 366 days. The value 366 serves as a unique token indicating that the time series is predicted not to experience any shortfalls within an annual fiscal year.
Notationally, our data consists of $(d_{i},X_{i},t_{i},y_{i}),\;i=1,2,\ldots,n$, where $X_{i}$ represents the $28$-day sequence features/covariates, $t_{i}$ is the time in days until a shortfall, $y_{i}$ is an indicator with one signifying that the time series resulted in a shortfall and 0 otherwise, and $d_{i}$ is a tuple containing three identifiers for the site, Ford plant, and vehicle part. Figure 1 illustrates the Pearson correlation among $21$ features, revealing no pairs of features with perfect correlation. Furthermore, Figure 2 demonstrates that roughly $25$% of all the time series are right-censored.
@@ 4 Methodology
This section details our methodology for modeling the dataset discussed previously, to predict the time until shortfall.
Given a sample $(d_{1},X_{1},t_{1},y_{1}),(d_{2},X_{2},t_{2},y_{2}),\ldots,(d_{n},X_{n},t_{n},y_{n})$, we model the times until disruption using a Survival Analysis model, as described in Section 2.1. Specifically, our chosen hazard function has a logistic form but with a non-linear term $\phi_{k}$. The function $\phi_{k}(\cdot),k=1,2,\ldots,366$ is a smooth, nonlinear function that models the relationship between the feature $X_{i}$ and the time until disruption $t_{i}$.
$h(\tau_{k}|X_{i})=\frac{1}{1+\exp(-\phi_{k}(X_{i}))},\;k=1,2,\ldots,366$ (5)
We estimate the functions $\phi_{1},\ldots,\phi_{366}$ by a Deep Learning Neural Network, Seq2Seq with bidirectional GRUs and Attention mechanism described in Section 2.3. The feature $X_{i}$ represents a sequence of 21 metrics collected over 28 days, as detailed in Section 3. This sequence is fed into the Encoder, which transforms it into 28 hidden states and a context vector—a weighted average of these hidden states with Attention Mechanism. The Encoder then processes these hidden states and the context vector to generate 366 elements, $\phi_{1},\phi_{2},\ldots,\phi_{366}$. These values, $\{\phi_{k}\}_{k=1}^{366}$, are used to compute the hazard functions, which are crucial to the final loss function employed to train the neural network via backpropagation. The loss function is the negative log-likelihood (3), given by:
@@
In terms of model architecture, we propose a important change to the Seq2Surv model (li2022), specifically enhancing the Encoder Network to better model heterogeneity among plants, sites, and parts. By incorporating Neural Embeddings for plant, site, and part IDs, we aim to effectively capture the dataset’s diversity. We refer to our modified model as Heterogeneous Sequence-to-Survival.
As illustrated in Figure 3 each site ID, plant ID, and vehicle part ID is transformed into vector embeddings using neural network embedding layers. These embeddings are then concatenated and passed through two fully connected neural network layers, which generate non-linear group effect signals. These signals are subsequently added to the hidden states of the bi-directional GRUs and concatenated with the context vector. This methodological adjustment is crucial for accounting for the heterogeneity in the dataset. Empirical observations indicate that the disruption times for different combinations of sites, plants, and parts can exhibit distinct and sometimes contradictory behaviors, highlighting the need for our approach to capture these nuanced variations. The Encoder encapsulates the essential representation for each $28$-day sequence feature $X_{i}$. Intuitively, this representation consists of two components: an individual-specific component and a group component related to the combination of site, plant, and part values. The individual-specific information is directly inferred from the feature values of $X_{i}$, ensuring that the unique characteristics of each sequence are accurately modeled. Meanwhile, the group component is derived from the site, plant, and part values, allowing the model to incorporate the distinct group behaviors from the data. The individual representation captures the idiosyncrasies of each sequence, while the group representation integrates the broader, systematic effects associated with different combinations of site, plant, and part values. This modified architecture enhances the model’s ability to predict shortfall times with greater accuracy and reliability, accommodating the inherent variability within the dataset.
@@ 5.1 Predictive Performance
North American Plant A 0.85 0.82 0.23 0.05 0.04 0.68
North American Plant B 0.95 0.9 0.22 0.02 0.00 0.75
North American Plant C 0.95 0.81 0.15 0.04 0.01 0.8
North American Plant D 0.95 0.86 0.21 0.03 0.01 0.75
North American Plant E 0.85 0.81 0.25 0.06 0.04 065
@@
In this section, we detail the approach for evaluating the proposed model at Ford to ensure its quality. The trained model estimates the time until a shortfall event occurs related to the shipment of vehicle parts between a supplier and a Ford plant. To assess the model’s performance, we modify and adapt the standard classification metrics to fit our specific setting.
In binary classification, the two standard metrics for evaluation are Precision and Recall (see buckland1994). Binary classification involves two labels: positive and negative. For a given classifier, True Positives (TP) are the instances where the classifier correctly identifies positives. True Negatives (TN) are the instances where the classifier correctly identifies negatives. False Positives are instances where the classifier incorrectly predicts positives when the true labels are negative. Conversely, False Negatives (FN) are instances where the classifier incorrectly predicts negatives when the true labels are positive. Formally, Precision and Recall are defined as,
@@
Precision and recall range from 00 to $1$. The closer these values are to $1$, the better the classifier’s performance. Specifically, recall quantifies the proportion of actual positive data points that the classifier successfully identifies. Precision, on the other hand, measures the proportion of predicted positive cases that are, in fact, positive. Informally, recall measures how effectively a classifier identifies positive events (shortfalls in this context), while precision measures the frequency with which the classifier avoids generating false alarms. These two metrics offer a comprehensive view of a model’s classification performance.
Given that the model proposed in Section 4 estimates the time until a shortfall, we first define a forecasting horizon, $\Delta$ (in days), and a margin-of-error, $\epsilon$ (in days), to evaluate its performance. We introduce the following definitions to adapt the concepts of True Positives, False Positives, True Negatives, and False Negatives to our specific problem. Given an observation time $t_{0}$, and the forecasting window $\Delta$ and margin-of-error $\epsilon$, $t_{0}+\Delta$ represents a time point $\Delta$ days from $t_{0}$. Adapted True Positives (ATP) are defined as the number of shortfalls that occur exactly $\Delta$ days from $t_{0}$ and are estimated by the model within $\epsilon$ days of $t_{0}+\Delta$. The difference between the model’s estimated time and $t_{0}+\Delta$ must fall within the range $[-\epsilon,\epsilon]$ days. Adapted False Positives (AFP) are instances where the model predicts a shortfall time exactly $\Delta$ days from $t_{0}$, but no shortfalls actually occur within $\epsilon$ days of $t_{0}+\Delta$.
Conversely, Adapted True Negatives (ATN) are instances where the model estimates the shortfall time to be more than $\Delta$ days from $t_{0}$, and no disruptions occur between $t_{0}$ and $t_{0}+\Delta$. Adapted False Negatives (AFN) are instances where the model predicts a shortfall time beyond $\Delta$ days from $t_{0}$, while shortfalls actually occur within $\Delta$ days from $t_{0}$.
With these definitions, Adapted Recall is computed as the ratio of Adapted True Positives (ATP) to the total of ATP and Adapted False Negatives (AFN). Likewise, Adapted Precision is the proportion of ATP relative to the sum of ATP and Adapted False Positives (AFP). Table 2 shows the metric performance for our proposed model. Additionally, we include the Normalized Confusion Matrix in Table 2. The Confusion Matrix offers a detailed breakdown of the four cases: True Positives, False Negatives, False Positives, and True Negatives. The normalization is done by dividing each case by the total number of cases.
@@
Seq2Surv on North American Plant A’s Data 0.62 0.3
Seq2Surv on North American Plant E’s Data 0.52 0.21
Heterogeneous Seq2Surv on North American Plant A’s Data 0.85 0.82
Heterogeneous Seq2Surv on North American Plan E’s Data 0.85 0.81
Table 3: Model Performance Comparison of Seq2Surv and Heterogeneous Seq2Surv on North American Plant A and B. Heterogeneity presents a critical challenge for Seq2Surv.
We conducted 20 iterations of Quality Assurance (QA) testing from August 1, 2023, to January 13, 2024 for the result in Table 2. For each observation week, we trained the model on the cumulative data up to that week and used it to predict disruptions in the near future. We computed Adapted Precision, Adapted Recall, and the Normalized Confusion Matrix using the above definitions for Adapted True Positives, True Negatives, False Positives, and False Negatives, with parameters forecasting horizon $\Delta=28$ days and margin-of-error $\epsilon=7$ days. The metrics in Table 2 are averaged over the 20 iterations. The Quality Assurance (QA) evaluation was conducted at five selected Ford manufacturing facilities in North America.
The metrics provided are the averages obtained from 20 iterations of QA testing. The model demonstrated on average precision exceeding $0.85$ and a recall surpassing $0.8$ across the five plants during the 20-week QA period. Additionally, the normalized confusion matrix indicates that the combined error rate of false positive and false negative errors was below $10$% on average. Furthermore, we present performance comparison between Seq2Surv, as described in li2022, and our Heterogeneous Seq2Surv proposed in Section 4, as shown in Table 3. The heterogeneity in plant, supplier, and part groups poses significant challenges for Seq2Surv. However, by incorporating heterogeneous behaviors in the data into the model, our Heterogeneous Seq2Surv demonstrates substantially improved performance. For a comparison between Seq2Surv and other models, please refer to li2022.
@@ 6 Conclusion
In this paper, we first outline the process of constructing a dataset of multivariate time series for forecasting first-tier supply chain disruptions. We carefully selected features representing key aspects of classical Factory Physics, such as capacity, variability, utilization, and processing time.
The dataset, comprising over five hundred thousand individual time series, presents significant technical challenges due to its complexity and scale. Although these time series exhibit some commonalities, they also demonstrate substantial heterogeneity within specific subgroups, making traditional industrial and statistical models inadequate. To address these challenges, we propose a novel methodology that integrates an enhanced Attention Sequence-to-Sequence Deep Learning architecture with Neural Network Embeddings to model group effects, combined with a Survival Analysis model. This approach is designed to effectively capture complex heterogeneous data patterns related to operational disruptions. Our model has achieved strong performance, with at least $85$% precision and $80$% recall during the Quality Assurance (QA) phase across Ford’s five North American plants.
To mitigate the criticism of Deep Learning models as “black boxes,” we employed the SHAP (SHapley Additive exPlanations) framework to elucidate feature importance in the model’s predictions. This technique provides valuable insights that can guide actionable strategies for business partners. Our work highlights the potential of advanced machine learning techniques to manage and mitigate supply chain risks effectively within the automotive industry.